# Fenchel-Young Variational Inference

### v2

@@ Appendix A $\rho$-MaxEnt and relation with Amari’s $\alpha$- and $q$-divergences
There are two relevant approaches to derive the same $\rho$-exponential family: maximize the $(2-\rho)$-entropy subject to linear expectation constraints, or maximize the $\rho$-entropy subject to escort expectation constraints. Amari and Ohara [4] follow the latter approach and are interested in cases where the two densities $q,p$ belong to the $\rho$-expfam that maximizes the Tsallis $\rho$-entropy subject to escort distribution constraints. In this paper, we focus on the $(2-\rho)$-exponential family of distributions that arises from the Tsallis $\rho$-MaxEnt principle subject to standard linear expectation constraints, following Naudts [43] and Martins et al. [39]. These densities are the $\Omega_{\rho}$-regularized prediction maps $p(z;\eta)=\exp_{2-\rho}[\eta z-A_{\rho}(\eta)]$ (c.f. Martins et al. [39, Def. 2]). Overloading $\eta$ to denote both a function and canonical parameters, the scoring function is $\eta(z)=\eta z$. The FY loss in Eq. (6) above, induced by the negentropy $\Omega$ essentially compares two members of the $(2-\rho)$-exponential family: $L_{\Omega}(\eta;q)=\Omega_{\rho}(q)-\Omega_{\rho}(p(z;\eta))-\eta^{\textnormal{T}}(\mathbb{E}_{p_{\eta}}[z]-\mathbb{E}_{q}[z])$ (c.f. Martins et al. [39, Eq. 14]).
Starting with regularized prediction maps like Martins et al. [39, Def. 2] the index $\rho$ ($\alpha$ in their notation) of the Tsallis $\rho$-entropy is specified first, and then the associated deformed family is defined using $\exp_{2-\rho}$, based on the $\rho$-MaxEnt problem under standard constraints [39, Eq. 10]. On the other hand, if the index $\rho$ is used to write a $\rho$-exponential family first, then the associated Tsallis entropy that is maximized, again with standard constraints, is the $(2-\rho)$-entropy (see also Naudts [44, Chapters 7, 8] and Wada and Scarfone [55], Wada et al. [56].) Amari and Ohara [4] use the index $\rho$ to write the $\rho$-deformed family using $\exp_{\rho}$. They concentrate on the “integration constant” that they call the $\rho$-log partition function (their $\psi$ is the same as the $A_{\alpha}$ of Martins et al. [39], given correct mapping of indices; we use $A_{\rho}$ above). Then, Amari and Ohara [4] derive the dual of $\psi$ which they call $\rho$-negentropy. Their $\rho$-divergence is the Bregman divergence based on the $\rho$-log partition, which is a useful quantity. Martins et al. [39, Prop. 10] also defines useful quantities: they derive the dual of the Tsallis negentropy to use in the FY loss. Suppose that we use $\rho$ as Amari and Ohara [4] for the $\rho$-exponential. Then Martins et al. [39, Prop. 9] describes how to take the associated $(2-\rho)$-Tsallis negentropy as $\Omega$ and derive its dual $\Omega^{*}$, which is not the log normalizer $A_{\rho}$.
@@
Let $\mathcal{Q}$ be a convex set of probability distributions over a space $\mathcal{Z}$, $\ell(x;z)$ a measurable loss function, and $\eta:\mathcal{Z}\rightarrow\mathbb{R}$ a scoring function (possibly a log-prior). Let $\Omega:\mathcal{Q}\to\mathbb{R}\cup\{+\infty\}$ be a proper, lower semi-continuous, convex regularizer that induces a Fenchel-Young loss:
$L_{\Omega}(\eta;q):=\Omega(q)+\Omega^{*}(\eta)-\mathbb{E}_{Z\sim q}[\eta(Z)].$
@@
$q_{x}^{\star}=\arg\min_{q\in\mathcal{Q}}\mathbb{E}_{Z\sim q}[\ell(x;Z)]+L_{\Omega}(\eta;q)$ (24)
@@
$=\exp_{2-\rho}\left[1+\eta(z)-\ell(x;z)-A_{\rho}\right],$ (26)
@@
$q_{x}^{\star}=\arg\min_{q\in\mathcal{Q}}\mathbb{E}_{Z\sim q}\left[\ell(x;z)\right]+L_{\Omega}(\eta;q)$
@@
that by definition maximizes the $\rho$-Tsallis entropy subject to an expectation constraint. Using Martins et al. [39, Prop. 8] we can express $q_{x}^{\star}(z)$ as a member of the $(2-\rho)$-exponential family. ∎
@@ Appendix B Clustering with FY EM
The FY EM algorithm described in $\S$3.3 recovers standard EM for GMMs when:
  • $x_{1},...,x_{N}\in\mathbb{R}^{d}$ are $d$-variate observations and the latent variables $z_{1},...z_{N}\in\mathcal{Z}=\{1,...,K\}$ indicate to which component each observation $x_{i}$ belongs;
  • $q_{i}=\bigl{(}q_{i}(1),...,q_{i}(K)\bigr{)}\in\!\Delta(\mathcal{Z})$ is the variational approximation of the distribution of $z_{i}$;
  • $\eta\in\mathbb{R}^{K}$ contains the logarithms of the prior probabilities of the $K$ components, denoted $\pi_{1},...,\pi_{K}$, i.e., $\eta_{1}=\log\pi_{1},...,\eta_{K}=\log\pi_{K}$;
  • $\ell(x_{i},z,\theta)=-\log\mathcal{N}(x_{i};\mu_{z},\Sigma_{z})$, for z=1,,Kz=1,...,K, where $\mu_{z}$ and $\Sigma_{z}$ are, respectively, the mean and covariance of component $z$; and
  • $\Omega(q)=-H(q)$, thus $\nabla\Omega^{*}=$ softmax.
@@
$\hat{q}_{i}^{\mbox{\scriptsize new}}(z)=\mbox{softmax}\bigl{(}\eta_{z}+\log\mathcal{N}(x_{i};\hat{\mu}_{z},\hat{\Sigma}_{z})\bigr{)}=\pi_{z}\;\mathcal{N}(x_{i};\hat{\mu}_{z},\hat{\Sigma}_{z})\biggl{(}\sum_{j=1}^{K}\pi_{j}\;\mathcal{N}(x_{i};\hat{\mu}_{j},\hat{\Sigma}_{j})\biggr{)}^{-1}\!,$
for i=1,,N,i=1,...,N, z=1,,Kz=1,...,K, where $\hat{\mu}_{z}$, $\hat{\Sigma}_{z}$ are the current parameter estimates. This is the E-step of standard EM for GMM estimation [23]. The M-step corresponds to maximum weighted log-likelihood estimation, yielding, for $z_{i}=1,...,K$,
$\hat{\mu}_{z}^{\mbox{\scriptsize new}}$ $=\frac{\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{\scriptsize new}}(z)\;x_{i}}{\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{\scriptsize new}}(z)},\quad\hat{\Sigma}_{z}^{\mbox{\scriptsize new}}=\frac{\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{\scriptsize new}}(z)\;(x_{i}-\hat{\mu}_{z}^{\mbox{\scriptsize new}})(x_{i}-\hat{\mu}_{z}^{\mbox{\scriptsize new}})^{\top}}{\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{\scriptsize new}}(z)},$
$\hat{\eta}_{z}^{\mbox{\scriptsize new}}$ $=\log\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{\scriptsize new}}(z)-\log N.$
@@
$\hat{q}_{i}^{\mbox{\scriptsize new}}\!=\mbox{argmax}\bigl{(}\pi_{1}\;\mathcal{N}(x_{i};\hat{\mu}_{1},\hat{\Sigma}_{1}),...,\pi_{K}\;\mathcal{N}(x_{i};\hat{\mu}_{K},\hat{\Sigma}_{K})\bigr{)},$
where $\mbox{argmax}(u_{1},...,u_{K})$ of a tuple of $K$ numbers returns a vector with $1/m$ in the entries with the same indices as the $m$ maxima and zero everywhere else. If the maximum is unique, this is just the one-hot representation of the maximizer. This hard version of EM coincides with the classification EM (CEM) algorithm proposed by Celeux and Govaert [13]. The M-step is the same as in standard EM.
Experimental Details.
@@
We evaluate the quality of unsupervised clustering of each algorithm using three metrics: adjusted mutual information [54], adjusted Rand index [48], and silhouette score [51]. Each model is trained and evaluated across five different random seeds, and the average and standard deviation of the metrics are reported in Table 1, showing that sparse EM outperforms standard and hard EM in two of three standard metrics. See also Figure 1 for a qualitative illustration. In Figure 2 we vary $\rho=0.1,0.5,0.9,1.0,1.1,1.5,2.0,3.0$ and evaluate the same three metrics when clustering with the corresponding EM. In the right most image of Figure 2 we verify that the sparsity of the E-step corresponds to the value of $\rho$ for the same values of $\rho$. We quantify the E-step sparsity by counting the number of clusters assigned 0 mixing proportion at the last iteration of each EM run and averaging across samples. Since the total number of clusters is four, the greatest number of components assigned 0 proportion is three.
Appendix C MNIST and FashionMNIST
@@
$\sum\limits_{i=1}^{N}\underset{Z\sim q_{i}}{\mathbb{E}}\bigl{[}\log p_{\theta}(x_{i}|Z)\bigr{]}-\beta D_{\text{\small KL}}(q_{i}\|p_{Z}),$ (27)
where $q_{i}=\mathcal{N}\bigl{(}\mu_{\phi}(x_{i}),\mbox{diag}\bigl{(}\sigma^{2}_{\phi}(x_{i})\bigr{)}\bigr{)}$ and
$p_{\theta}(x_{i}|z_{i})\sim\textnormal{Bernoulli}(\sigma(f_{\theta}(z_{i}));$
$\sigma$ denotes the sigmoid function, and $f_{\theta}$ a neurally parametrized function.
Experimental Details.
@@
We report the $\ell_{1}$ reconstruction error in Table 2, i.e. the absolute distance between the model’s reconstructed samples and the original input images, evaluated on the test split for both MNIST and FashionMNIST datasets. [^6] MNIST is Public Domain, and FashionMNIST is under an MIT License [58]. The reconstruction error is the only standard and interpretable metric that is comparable across the VAE alternatives we consider. The FY ELBO has a log scale only in the standard VAE case, which is lost when either the observation or latent variable model is based on a Tsallis entropy with $\rho>1$. The first row of Table 2 corresponds to the standard VAE with a Gaussian latent variable. The Biweight and Epanechnikov densities are limited support distributions of the latent variable, with Tsallis entropies corresponding to $\rho=3/2$ $\rho=2$ respectively. The best reconstruction errors are obtained with latent $(2-\rho)$-Gaussian encoders combined with 2-entmax observation models.
Appendix D Document Fenchel-Young $\beta$-VAE
@@
Similarly with one of the cases we compared in the image VAE, namely the sparsemax observation model combined with a Gaussian latent variable in the previous section, we experiment with a FY $\beta$-VAE applied to unsupervised word embeddings to directly model the sparsity of the empirical distribution. In this experiment we focus on tuning the level of sparsity in the observation model rather than using only $\rho^{\prime}=2$. The baseline neural variational document model (NVDM) proposed by Miao et al. [40] is a standard VAE (with prior $p_{Z}=\mathcal{N}(0,I)$): it models the joint probability of observed word counts (or proportions) in a document $x_{i}$, i=1,,Ni=1,...,N, and of continuous latent semantic embeddings $z_{i}$, by optimizing the same objective as (27) and replacing the multivariate Bernoulli with a multinomial observation model,
$p_{\theta}(x_{i}|z_{i})\sim\textnormal{Multinomial}(\textnormal{softmax}(Wz_{i})),$
where $\theta=W\in\mathbb{R}^{|D|\times K}$, $K$ is the dimension of the latent multivariate Gaussian, and $|D|$ is the size of the observed vocabulary. The model parameters are learned by stochastic gradient backpropagation.
Our FY $\beta$-VAE model uses the same multivariate Gaussian latent $q$ and prior $p_{Z}$, combined with a FY observation model $L_{\Psi}(\theta(z);x)$:
$J(\phi,\theta)=\sum\limits_{i=1}^{N}\underset{Z\sim q}{\mathbb{E}}\left[L_{\Psi}(\theta(Z);x_{i})\right]+\beta D_{\text{\footnotesize KL}}(q_{i}\|p_{Z}),$
@@
where $\Psi$ is a Tsallis $\rho^{\prime}$-negentropy, dom($\Psi)=\Delta_{|D|-1}$, and $\rho^{\prime}$-entmax replaces softmax in the $L_{\Psi}$ decoder. Unlike the standard VAE formulation, the deformed exponential family distribution parametrized by $\theta(z)$ can assign zero probability to vocabulary categories that do not occur in a given document, directly modeling the sparse empirical distribution.
@@
Experimental Details. We use the 20NewsGroups [^7] http://qwone.com/~jason/20Newsgroups/, [35]. dataset, which consists of 11,314 training samples and 7,531 test samples (documents, represented as bags of words). We assume a decoder parameterizing a sparse discrete distribution over the vocabulary generates all the words in a document independently, given a continuous Gaussian latent variable $Z$, i.e. a dense, continuous document representation. Following Miao et al. [40], we assume that each dimension in the latent space represents a “topic”, which is a simplification of the classical discrete latent variable topic model of [8].
@@
The first 1,000 test samples were used as the validation set for hyperparameter tuning. We perform a search over entmax $\rho^{\prime}$, which controls the decoder’s sparsity, learning rate, batch size, number of epochs, and the number of topics using Optuna [1] across 100 trials for both the baseline and FY NVDM models. Finally, we rerun the experiment with the best hyperparameters using five different random seeds and report the reconstruction error (RE) defined by the $\ell_{1}$ distance between the empirical data distribution and the learned observation model distribution. For the baseline NVDM model over 100 Optuna trials the hyperparameters that yield the best reconstruction error of 1.484 reported in Table 3 are batch size = 64, number of alternating epochs updating the decoder and encoder parameters = 1. For the Fenchel-Young $\beta$-VAE over 100 Optuna trials the hyperparameters that yield the best reconstruction error of 1.384 reported in Table 3 are $\rho^{\prime}=1.314$, batch size = 1024, number of alternating epochs updating the decoder and encoder parameters = 3. In both cases $\beta=0.01$, learning rate $5\times 10^{-5}$, number of training epochs = 500, and dimension of the latent Gaussian = 50. Similarly to the image VAE experiments we use $\beta=0.01$ ensures that the regulazation term is non trivially small. Both models converge based on training objective curves, and in preliminary experiments we empirically verified that for $\rho^{\prime}$ close to 1, training and validation behavior was similar to the NVDM baseline. Again the objectives are are not numerically comparable to each other, nor is expected $\Psi_{r}^{\prime}$ term with the log scale regularization term. Again $\ell_{1}$ reconstruction error is a metric that is comparable between the NVDM and the FY NVDM.
Our sparse support observation model FYVAE outperforms the NVDM baseline method [40] by 10.68% in absolute reconstruction error, as reported in Table 3.
@@
NVDM 1.4984 ± 0.0009
FY VAE 1.3384 ± 0.0008
@@
Figure 3 shows average coherence scores, measured by the pairwise similarity between embedding representations of the top-$k$ words for each latent “topic” dimension, using a sentence-transformer model [49]. [^8] We use the all-mpnet-base-v2 model, which is available at huggingface.co/sentence-transformers/all-mpnet-base-v2, under Apache-2.0. FYVAE consistently achieves better or similar coherence scores across varying values of $k$ (number of words) than the NVDM model, confirming that the latent document representations learned by FYVAE are more semantically coherent and better aligned with meaningful word groupings.

### v3

@@ Appendix A RELATION BETWEEN FY EXPRESSIONS, $\alpha$- AND $q$-DIVERGENCES
There are two alternative approaches to derive the same $\rho$-exponential family: Maximize the $(2-\rho)$-entropy subject to linear expectation constraints, or maximize the $\rho$-entropy subject to escort expectation constraints. amari2011geometry follow the latter approach and are interested in characterizing the divergence between densities $q,p$ that belong to the $\rho$-expfam obtained by maximizing the Tsallis $\rho$-entropy subject to escort distribution constraints. In this paper, we focus on the $(2-\rho)$-exponential family of distributions that arises from the Tsallis $\rho$-MaxEnt principle subject to standard linear expectation constraints, following naudts2009q and martins2022sparse. These densities are the $\Omega_{\rho}$-regularized prediction maps $p(z;\eta)=\exp_{2-\rho}[\eta^{\top}z-A_{\rho}(\eta)]$ ([martins2022sparse, Def. 2]). Overloading $\eta$ to denote both a function and canonical parameters, the scoring function is $\eta(z)=\eta^{\top}z$. The FY expression in Eq. (15) above, induced by the negentropy $\Omega$ compares two members of the $(2-\rho)$-exponential family: $D_{\Omega}(\eta;q)=\Omega_{\rho}(q)-\Omega_{\rho}(p(z;\eta))-\eta^{\top}(\mathbb{E}_{p_{\eta}}[z]-\mathbb{E}_{q}[z])$ ([martins2022sparse, Eq. 14]).
Starting with regularized prediction maps like [martins2022sparse, Def. 2] the index $\rho$ of the Tsallis entropy is specified first, and then the associated deformed family is defined using $\exp_{2-\rho}$, based on the $\rho$-MaxEnt problem under standard constraints [martins2022sparse, Eq. 10]. On the other hand, if the index $\rho$ is used to write a $\rho$-exponential family first, then the associated Tsallis entropy that is maximized, again with standard constraints, is the $(2-\rho)$-entropy (see Chapters 7 and 8 in naudts2011generalised and wada2005connections, wada2012relationships).
@@
Let $\mathcal{Q}$ be a set of probability distributions over a space $\mathcal{Z}$, let $\ell(x;z)$ be a measurable loss function, and let $\eta:\mathcal{Z}\rightarrow\mathbb{R}$ be a scoring function (possibly a log-prior). $\Omega:\mathcal{Q}\to\mathbb{R}\cup\{+\infty\}$ be a proper, lower semi-continuous, convex regularizer that induces a Fenchel-Young discrepancy:
$D_{\Omega}(\eta;q):=\Omega(q)+\Omega^{*}(\eta)-\mathbb{E}_{z\sim q}[\eta(z)].$
@@
$q_{x}^{\star}=\arg\min_{q\in\mathcal{Q}}\penalty 10000\ \mathbb{E}_{z\sim q}[\ell(x;z)]+L_{\Omega}(\eta;q)$
@@
$q_{x}^{\star}(z)$ $=\left[1+(1-\rho)\left(\eta(z)-\ell(x;z)-A_{\rho}\right)\right]_{+}^{\frac{1}{\rho-1}}$
$=\exp_{2-\rho}\left[1+\eta(z)-\ell(x;z)-A_{\rho}\right],$
@@
$q_{x}^{\star}=\arg\min_{q\in\mathcal{Q}}\mathbb{E}_{z\sim q}\left[\ell(x;z)\right]+D_{\Omega}(\eta;q)$
@@
that maximizes the $\rho$-Tsallis entropy subject to an expectation constraint [martins2022sparse, App. C.4]. Using [martins2022sparse, Prop. 8] we can express $q_{x}^{\star}(z)$ as a member of the $(2-\rho)$-exponential family. ∎
@@ Appendix B CLUSTERING WITH FY EM
The FY EM algorithm described in $\S$3.3 recovers standard EM for GMMs when: $x_{1},...,x_{N}\in\mathbb{R}^{d}$ are $d$-variate observations and the latent variables $z_{1},...z_{N}\in\mathcal{Z}=\{1,...,K\}$ indicate to which component each observation $x_{i}$ belongs; $q_{i}=\bigl(q_{i}(1),...,q_{i}(K)\bigr)\in\!\Delta(\mathcal{Z})$ is the variational approximation of the distribution of $z_{i}$; $\eta\in\mathbb{R}^{K}$ contains the logarithms of the prior probabilities of the $K$ components, denoted $\pi_{1},...,\pi_{K}$, i.e., $\eta_{1}=\log\pi_{1},...,\eta_{K}=\log\pi_{K}$; $\ell(x_{i},z,\theta)=-\log\mathcal{N}(x_{i};\mu_{z},\Sigma_{z})$, for z=1,,Kz=1,...,K, where $\mu_{z}$ and $\Sigma_{z}$ are, respectively, the mean and covariance of component $z$; and $\Omega(q)=-H(q)$, thus $\nabla\Omega^{*}=$ softmax.
@@
$\hat{q}_{i}^{\mbox{new}}(z)=\mbox{softmax}\bigl(\eta_{z}+\log\mathcal{N}(x_{i};\hat{\mu}_{z},\hat{\Sigma}_{z})\bigr)=\pi_{z}\;\mathcal{N}(x_{i};\hat{\mu}_{z},\hat{\Sigma}_{z})\biggl(\sum_{j=1}^{K}\pi_{j}\;\mathcal{N}(x_{i};\hat{\mu}_{j},\hat{\Sigma}_{j})\biggr)^{-1}\!,$
for i=1,,N,i=1,...,N, z=1,,Kz=1,...,K, where $\hat{\mu}_{z}$, $\hat{\Sigma}_{z}$ are the current parameter estimates. This is the E-step of standard EM for GMM estimation [Figueiredo2002]. The M-step corresponds to maximum weighted log-likelihood estimation, yielding, for $z_{i}=1,...,K$,
$\hat{\mu}_{z}^{\mbox{new}}$ $=\frac{\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{new}}(z)\;x_{i}}{\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{new}}(z)},\quad\hat{\Sigma}_{z}^{\mbox{new}}=\frac{\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{new}}(z)\;(x_{i}-\hat{\mu}_{z}^{\mbox{new}})(x_{i}-\hat{\mu}_{z}^{\mbox{new}})^{\top}}{\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{new}}(z)},\hat{\eta}_{z}^{\mbox{new}}$ $=\log\sum\limits_{i=1}^{N}\hat{q}_{i}^{\mbox{new}}(z)-\log N.$
@@
$\hat{q}_{i}^{\mbox{new}}\!=\mbox{argmax}\bigl(\pi_{1}\;\mathcal{N}(x_{i};\hat{\mu}_{1},\hat{\Sigma}_{1}),...,\pi_{K}\;\mathcal{N}(x_{i};\hat{\mu}_{K},\hat{\Sigma}_{K})\bigr),$
where $\mbox{argmax}(u_{1},...,u_{K})$ of a tuple of $K$ numbers returns a vector with $1/m$ in the entries with the same indices as the $m$ maxima and zero everywhere else. If the maximum is unique, this is just the one-hot representation of the maximizer. This hard version of EM coincides with the classification EM (CEM) algorithm proposed by Celeux. The M-step is the same as in standard EM.
Experiment Details.
@@
We evaluate the quality of unsupervised clustering of each algorithm using three metrics: adjusted mutual information [Vinh], adjusted Rand index [Rand], and silhouette score [Rousseeuw]. Each model is trained and evaluated across five different random seeds, and the average and standard deviation of the metrics are reported in Table 1, showing that sparse EM outperforms standard and hard EM in two of three standard metrics. See also Figure 2 for a qualitative illustration. In Figure 3 we vary $\rho=0.1,0.5,0.9,1.0,1.1,1.5,2.0,3.0$ and evaluate the same three metrics when clustering with the corresponding EM. In the right most image of Figure 3 we verify that the sparsity of the E-step corresponds to the value of $\rho$ for the same values of $\rho$. We quantify the E-step sparsity by counting the number of clusters assigned zero mixing proportion at the last iteration of each EM run and averaging across samples. Since the total number of clusters is four, the greatest number of components assigned zero proportion is three.
Appendix C MNIST AND FMNIST FY VAES
@@
$\sum\limits_{i=1}^{N}\underset{z\sim q_{i}}{\mathbb{E}}\bigl[\log p_{\theta}(x_{i}|z)\bigr]-\beta D_{\text{KL}}(q_{i}\|p_{Z}),$ (32)
where $q_{i}=\mathcal{N}\bigl(\mu_{\phi}(x_{i}),\mbox{diag}\bigl(\sigma^{2}_{\phi}(x_{i})\bigr)\bigr)$ and $p_{\theta}(x_{i}|z_{i})\sim\textnormal{Bernoulli}(\sigma(f_{\theta}(z_{i}))$; $\sigma$ denotes the sigmoid function, and $f_{\theta}$ a neurally parametrized function.
Experiment Details.
@@
We report the $\ell_{1}$ reconstruction error in Table 2, i.e. the absolute distance between the model’s reconstructed samples and the original input images, evaluated on the test split for both MNIST and FashionMNIST datasets. The reconstruction error is the only standard and interpretable metric that is comparable across the VAE alternatives we consider. The FY ELBO has a log scale only in the standard VAE case, which is lost when either the observation or latent variable model is based on a Tsallis entropy with $\rho>1$. The first row of Table 2 corresponds to the standard VAE with a Gaussian latent variable. The Biweight and Epanechnikov densities are limited support distributions of the latent variable, with Tsallis entropies corresponding to $\rho=3/2$ $\rho=2$ respectively. The best reconstruction errors are obtained with latent $(2-\rho)$-Gaussian encoders combined with 2-entmax observation models.
Appendix D DOCUMENT FY $\beta$-VAE
@@
Similarly with one of the cases we compared in the image VAE, namely the sparsemax observation model combined with a Gaussian latent variable in the previous section, we experiment with a FY $\beta$-VAE applied to unsupervised word embeddings to directly model the sparsity of the empirical distribution. In this experiment we focus on tuning the level of sparsity in the observation model rather than using only $\rho^{\prime}=2$. The baseline neural variational document model (NVDM) proposed by miao2016neural is a standard VAE (with prior $p_{Z}=\mathcal{N}(0,I)$): it models the joint probability of observed word counts (or proportions) in a document $x_{i}$, i=1,,Ni=1,...,N, and of continuous latent semantic embeddings $z_{i}$, by optimizing the same objective as (32) and replacing the multivariate Bernoulli with a Multinomial observation model, $p_{\theta}(x_{i}|z_{i})\sim\textnormal{Multinomial}(\textnormal{softmax}(Wz_{i}))$, where $\theta=W\in\mathbb{R}^{|D|\times K}$, $K$ is the dimension of the latent multivariate Gaussian, and $|D|$ is the size of the observed vocabulary. The model parameters are learned by stochastic gradient backpropagation.
@@
Our FY $\beta$-VAE model uses the same multivariate Gaussian latent $q$ and prior $p_{Z}$, combined with a FY observation model $D_{\Psi}(\theta(z);x)$: $J(\phi,\theta)=\sum\limits_{i=1}^{N}\underset{Z\sim q}{\mathbb{E}}\left[D_{\Psi}(\theta(Z);x_{i})\right]+\beta D_{\text{KL}}(q_{i}\|p_{Z})$, where $\Psi$ is a Tsallis $\rho^{\prime}$-negentropy, dom($\Psi)=\Delta_{|D|-1}$, and $\rho^{\prime}$-entmax replaces softmax in the $D_{\Psi}$ decoder. Unlike the standard VAE formulation, the deformed exponential family distribution parametrized by $\theta(z)$ can assign zero probability to vocabulary categories that do not occur in a given document, directly modeling the sparse empirical distribution.
@@
Experiment Details. We use the 20NewsGroups [^4] http://qwone.com/~jason/20Newsgroups/, [lang1995newsweeder]. dataset, which consists of 11,314 training samples and 7,531 test samples (documents, represented as bags of words). We assume a decoder parameterizing a sparse discrete distribution over the vocabulary generates all the words in a document independently, given a continuous Gaussian latent variable $Z$, i.e. a dense, continuous document representation. Following miao2016neural, we assume that each dimension in the latent space represents a “topic”, which is a simplification of the classical discrete latent variable topic model of [blei2003latent].
@@
The first 1,000 test samples were used as the validation set for hyperparameter tuning. We perform a search over entmax $\rho^{\prime}$, which controls the decoder’s sparsity, learning rate, batch size, number of epochs, and the number of topics using Optuna [optuna2019] across 100 trials for both the baseline and FY NVDM models. Finally, we rerun the experiment with the best hyperparameters using five different random seeds and report the reconstruction error (RE) defined by the $\ell_{1}$ distance between the empirical data distribution and the learned observation model distribution. For the baseline NVDM model over 100 Optuna trials the hyperparameters that yield the best reconstruction error of 1.484 reported in Table 3 are batch size = 64, number of alternating epochs updating the decoder and encoder parameters = 1. For the Fenchel-Young $\beta$-VAE over 100 Optuna trials the hyperparameters that yield the best reconstruction error of 1.384 reported in Table 3 are $\rho^{\prime}=1.314$, batch size = 1024, number of alternating epochs updating the decoder and encoder parameters = 3. In both cases $\beta=0.01$, learning rate $5\times 10^{-5}$, number of training epochs = 500, and dimension of the latent Gaussian = 50. Similarly to the image VAE experiments we use $\beta=0.01$ ensures that the regularization term is non trivially small. Both models converge based on training objective curves, and in preliminary experiments we empirically verified that for $\rho^{\prime}$ close to 1, training and validation behavior was similar to the NVDM baseline. Again the objectives are are not numerically comparable to each other, nor is expected $\Psi_{r}^{\prime}$ term with the log scale regularization term. Again $\ell_{1}$ reconstruction error is a metric that is comparable between the NVDM and the FY NVDM.
@@
NVDM 1.4984 ± .0004
FY VAE 1.3384 ± .0004
Our sparse support observation model FYVAE outperforms the NVDM baseline method [miao2016neural] by 10.68% in absolute reconstruction error, as reported in Table 3. Figure 4 shows average coherence scores, measured by the pairwise similarity between embedding representations of the top-$k$ words for each latent “topic” dimension, using a sentence-transformer model [reimers-gurevych-2019-sentence]. [^5] We use the all-mpnet-base-v2 model, which is available at huggingface.co/sentence-transformers/all-mpnet-base-v2, under Apache-2.0. FYVAE consistently achieves better or similar coherence scores across varying values of $k$ (number of words) than the NVDM model, confirming that the latent document representations learned by FYVAE are more semantically coherent and better aligned with meaningful word groupings.